A disk between vertebrae in the spine is subjected to a shearing force of 530 N. Find its shear deformation taking it to have the shear modulus of 1 ? 109 N/m2. The disk is equivalent to a solid cylinder 0.700 cm high and 3.90 cm in diameter.
Answer #1

S = F/A/(?x/L)

=> .. ?x = F*L/(A*S)

?x =530*0.007/(3.14*0.0112*10^9)

=3.71/0.035*10^9

=106m

=0.106mm